# Improving absolute separability bounds for arbitrary dimensions

### v1

@@ 3 Improving sufficient criteria for absolute separability in arbitrary dimensions from convex geometry
Having set the main tools of our analysis, here we show how to improve present bounds of absolute separability based on convex geometrical methods. The set of states detected by the inverse of the reduction map (Theorem 3) for different values of $\alpha$ and the states enclosed in Gurvits-Barnum ball (Theorem 1) are not comparable. Nevertheless, since the set of AS states is convex and compact [32], merging both criteria with extreme values of $\alpha_{\pm}$ into a single convex hull can yield stronger separability conditions.
3.1 Detecting separability from few eigenvalues
@@
The subsets of AS states in $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ detected by Theorem 3, form simplexes in the space of eigenvalues. For the case of arbitrary dimensions $D=N\cdot M$, the vertices of these simplexes correspond to all permutations on the vector formed by the eigenvalues. Explicitly, for $\alpha=2$, the largest simplex corresponds to the vector $\left(\frac{1}{D+2},\cdots,\frac{1}{D+2},1-\frac{D-1}{D+2}\right)$ and its permutations, while for $\alpha=-1$, the vertices are settled by $\left(0,\frac{1}{D-1},\cdots,\frac{1}{D-1}\right)$ and its permutations. A schematic representation of these sets for $D=3$ is depicted in Figure 1 using barycentric coordinates $\{\tilde{\lambda}_{i}\}$, where the maximally mixed state (MMS) lies at the origin and pure states at the vertexes of the simplex of normalized states.
$\tilde{\lambda}_{1}$$\tilde{\lambda}_{2}$
Figure 1: Schematic representation of the set of states and absolutely separable states (AS) for the case of $D=3$ in barycentric coordinates. The normalized set of states is represented by the area enclosed in the blue vertices. The two simplexes obtained from Theorem 3, corresponding to the conditions $\lambda_{\min}(\rho)\geq 1/(D+2)$ (red dashed line) and the set with $\lambda_{\max}(\rho)\leq 1/(D-1)$ (orange dotted line), are also plotted. The thick green line displays the convex hull of both simplexes. Finally, the shaded zone fulfills the ordering $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}$ that we consider across this work without loss of generality. Note that the case $D=3$ does not correspond to a physical system, as it cannot factorize into any subsystems; however, it is depicted here to exemplify the general rule.
@@
The displayed simplexes of AS are, in fact, extreme due to the tightness of the values of $\alpha$ displayed in Eq. (3). Thus, the convex hull arising from the union of both simplexes will also be optimal, i.e., the biggest possible one. Through convexity of the AS set, it is also possible to upper bound it, since the vertex of a simplex with $\alpha\pm\varepsilon$ will not be AS. We formally state these facts in the following proposition.
Proposition 1.
@@
Let $\rho$ be a bipartite state acting in a Hilbert space $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ of dimension $D=N\cdot M$, and let $\{\lambda_{i}\}_{i=0}^{D-1}$ denote its corresponding eigenvalues in non-decreasing order $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,1,..,D-1}$. If $\rho$ is contained in the CH of the conditions presented in Theorem 3, then
$3\cdot\sum_{i=0}^{\lfloor\frac{D+1}{3}\rfloor-1}\lambda_{i}+\left(D+2-3\cdot\left\lfloor\frac{D+1}{3}\right\rfloor\right)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor}\geq 1$ (6)
@@
In Appendix A, we show how to numerically obtain the proposed solution as an ansatz given a specific dimension. In Appendix B, by assessing the Schur-convexity of the condition and majorization over the vectors of eigenvalues, we prove the statement. ∎
For low dimensional systems ($D\leq 16$), we verify that the converse relation also holds. Moreover, for higher dimensions, we obtain numerical evidence that only points contained in the CH are detected by the single inequality. Thus, Eq. 6 yields a proper criterion for AS.
@@
The condition provided by Eq. (6) holds for highly mixed states, close to the MMS, with the specified combination of the smallest eigenvalues large enough. Provided the ordering, $\{\lambda_{i+1}\geq\lambda_{i}\}$, one can replace $\lambda_{i+1}$ by $\lambda_{i}$ in Eq. (6) to yield new sufficient conditions for AS. Albeit being weaker, such conditions are based on fewer eigenvalues than the original theorem, which enables the detection of AS from incomplete knowledge of the spectrum. By iteratively applying the above substitution for $i\leq\kappa\leq\lfloor\frac{D+1}{3}\rfloor$, we derive a hierarchy of AS conditions as follows:
@@
$\kappa=0:$ $\quad 3\cdot\lambda_{0}+\cdots+3\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor-1}+(D+2-3\cdot\left\lfloor\frac{D+1}{3}\right\rfloor)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor}\geq 1,$ (7)
@@
$\kappa=1:$ $\quad 3\cdot\lambda_{0}+\cdots+(D+5-3\cdot\lfloor\frac{D+1}{3}\rfloor)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor-1}+\geq 1,$
@@
$\kappa=\lfloor\frac{D+1}{3}\rfloor:$ $\quad(D+2)\cdot\lambda_{0}\geq 1.$
Each level $\kappa$ in Eq. (7) is weaker than the previous $\kappa-1$, but also it does not require the knowledge of eigenvalue $\lambda_{\kappa}$. Notice that it also allows for some intermediate eigenvalues to be unknown. By considering the two smallest eigenvalues, we derive the weaker condition $(D-1)\lambda_{1}+3\lambda_{0}\geq 1$, which already includes the $2D$ vertices of both simplexes despite being suboptimal. The whole polytope of AS states is enclosed by all the possible permutations on the eigenvalues of inequality (6), but just as with the case of the simplexes in Figure 1, one linear inequality is enough for ordered eigenvalues (see Appendix A, B).
3.2 Improving Gurvits-Barnum criterion for arbitrary dimensions
Let us observe that the condition $\lambda_{\max}(\rho)\leq 1/(N\cdot M-1)$ is completely enclosed inside the Gurvits-Barnum ball. Thus, we focus on the set of states detected by the convex hull (CH) of the ball (Theorem 1) and the condition $\lambda_{\min}(\rho)\geq 1/(N\cdot M+2)$ (Theorem 3). This approach improves existing separability criteria for general dimension $D=N\cdot M$. Notably, the Hilbert-Schmidt distance between the vertices of the simplex and the maximally mixed state tends towards twice the radius of the ball as $N\cdot M\rightarrow\infty$. Figure 2 provides a 2D-sketch of the geometry of such a convex hull and how different regions are detected.
Figure 2: Detected absolutely separable (AS) states. Left panel: Schematic 2D representation of the convex hull (orange) of the Gurvits-Barnum ball (yellow) and our simplex (blue) in barycentric coordinates (c.f. Figure 1). Right panel: Purity $\mathrm{Tr}(\rho^{2})$ and minimal eigenvalue $\lambda_{\min}(\rho)$ of the detected states for a $2$-qutrit system. We distinguish the states detected as AS from Gurvits-Barnum ball (yellow region), the ones detected by the inverse of the reduction map (blue area), by both criteria (green region) and by its CH (orange area under the dotted curve). The points in the striped region do not correspond to physical states.
While an analytical expression for the CH (Figure 2) in higher dimensions is challenging to obtain, it is possible to state the problem as a disciplined convex program (DCP), as formalized in the next proposition.
@@ 3.2 Improving Gurvits-Barnum criterion for arbitrary dimensions
$\begin{array}[]{crl}\min_{\{\varrho_{1}\succeq 0,\varrho_{2}\succeq 0\}}&0&\\ \mbox{s.t.}&\rho&=\varrho_{1}+\varrho_{2}\\ &\lambda_{\min}(\varrho_{1})&\geq\frac{\mathrm{Tr}(\varrho_{1})}{D+2}\\ &\mathrm{Tr}(\varrho^{2}_{2})&\leq\frac{\mathrm{Tr}(\varrho_{2})^{2}}{D-1}\end{array}$ (8)
@@
If the problem is feasible, it implies that the algorithm indeed succeeded in finding $\{\varrho_{1},\varrho_{2}\}$ fulfilling the constraints. Hence, both states $\{\varrho_{1},\varrho_{2}\}$ are certified AS from Theorem 1 or Theorem 3, respectively. Finally, by convexity, their sum $\rho$ is also AS. ∎
@@ 3.3 Unifying separability criteria into a single convex program
Here, we outline how to generalize the disciplined convex program from Proposition 2 into potentially stronger sufficient conditions for separability. Although such criteria, as already discussed, arise from a plethora of different techniques, it is interesting to notice that all of them can be ultimately expressed as the collection of inequalities, $\{f_{i}(\rho)\leq 0\}$, where the functions $\{f_{i}\}$ are convex in the state $\rho$. Let us designate by $\mathcal{R}_{i}=\{\rho\succeq 0\mbox{ s.t. }f_{i}(\rho)\leq 0\}$ the set of (unnormalized) states detected by each criterion $i$. From the curvature of $\{f_{i}\}$, the sets $\{\mathcal{R}_{i}\}$ are convex as well. However, such sets, stemming from the difference conditions, may be incomparable. How can we build a new condition including them all? The following theorem offers a solution to this question and generalizes Proposition 2.
@@
Let $\rho$ be a quantum state, and $\{f_{i}(\rho)\leq 0\}$, with $\{f_{i}\}$ a set of convex functions leading to incomparable sets of states, each of them representing a sufficient conditions for separability. If the convex program
$\begin{array}[]{crl}\min_{\{\varrho_{i}\succeq 0\}}&0&\\ \mbox{s.t.}&\rho&=\sum_{i}\varrho_{i}\\ &\{f_{i}(\varrho_{i})&\leq 0\}\\ \end{array}$ (9)
@@
By construction, the previous theorem implicitly tests membership of the state $\rho$ in the convex hull [33] of the union of the detected sets $\bigcup_{i}\mathcal{R}_{i}:=\mathcal{R}$, in other words, it offers an answer to the question whether $\rho$ is contained in $\mathrm{CH}(\mathcal{R})$. Consequently, Theorem 4 potentially enables the certification of states $\rho$ that are not detected by any the criteria $\{f_{i}(\rho)\leq 0\}$ alone, but that nonetheless can be decomposed into states that are detected by such criteria. In turn, convex problems of the structure Eq. (9), are not only mathematically well-posed [34, 35], but for some instances they may admit polynomial-time solutions. In this regard, there is a number of available solvers tailored to these problems [36], as well as user-friendly parsers [37].

### v2

@@ 3 Improving sufficient criteria for absolute separability in arbitrary dimensions from convex geometry
As mentioned previously, the sets of states detected by the inverse of the reduction map, and those enclosed within the Gurvits-Barnum ball are not directly comparable. However, since the AS set is convex and compact [37], we can unify both criteria using convex geometrical tools into a single convex hull that yields stronger separability conditions.
3.1 Extremal points of the convex hull of AS criteria
@@
The set of AS states detected through Theorem 3, form simplexes in the space of eigenvalues. For arbitrary dimensions, $D=N\cdot M$, the vertices of these simplexes correspond to all permutations on the vector of eigenvalues. Explicitly, for $\alpha=2$, the largest simplex corresponds to the vector $\left(\frac{1}{D+2},\cdots,\frac{1}{D+2},1-\frac{D-1}{D+2}\right)$ and its permutations, while for $\alpha=-1$, the vertices are settled by $\left(0,\frac{1}{D-1},\cdots,\frac{1}{D-1}\right)$ and its permutations. A schematic representation of these sets for $D=3$ is depicted in Figure 1 using barycentric coordinates $\{\tilde{\lambda}_{i}\}$. In this representation, the maximally mixed state (MMS) lies at the origin and pure states correspond to the vertices of the simplex of normalized states.
$\lambda_{0}$$\lambda_{1}$$\lambda_{2}$
Figure 1: Schematic representation of quantum states and the AS set for $D=3$ in barycentric coordinates. Normalized states fill the area enclosed by the blue vertices. In red (dashed) the simplex corresponding to $\lambda_{\min}\geq 1/(D+2)$, in orange(dots) the one corresponding to $\lambda_{\max}\leq 1/(D-1)$ (see Theorem 3). In green the convex hull of both simplexes. The light shaded region fulfills the ordering $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}$ we consider. The dark shadow polytope highlights the intersection of the convex hull with the ordered zone. The figure is illustrative as $D=3$ does not correspond to any bipartite splitting.
@@
The displayed simplexes shown for AS are, in fact, extreme due to the tightness of the values of $\alpha$ obtained from Eq. (3). Thus, the convex hull arising from the union of both simplexes is optimal. The following Lemma fully describes, for given values of $\alpha_{\pm}$, the convex hull of the two simplexes.
Lemma 1.
@@
Let $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ be a normalized state in a Hilbert space of global dimension $D=N\cdot M$, and $\bm{\lambda}=\{\lambda_{i}\}_{i=0}^{D-1}$ the corresponding eigenvalues in a non-decreasing order, that is, $0\leq\lambda_{0}\leq\lambda_{1}\cdots\leq\lambda_{D-1}$, with $\sum_{i=0}^{D-1}\lambda_{i}=1$. Given $\alpha_{+}\geq 0$ and $\alpha_{-}\leq 0$, the vector $\bm{\lambda}$ is contained in the convex hull settled by conditions
$\lambda_{0}\geq\frac{1}{D+\alpha_{+}},$ (6)
$\lambda_{D-1}\leq\frac{1}{D+\alpha_{-}},$ (7)
if and only if
$K\cdot\sum_{i=0}^{c-1}\lambda_{i}+\left[D-K\cdot c+\alpha_{+}\right]\cdot\lambda_{c}\geq 1,$ (8)
where $K=\left(1-\frac{\alpha_{+}}{\alpha_{-}}\right)$, and $c=\left\lceil\frac{\alpha_{+}+\alpha_{-}(D-1+\alpha_{+})}{\alpha_{-}-\alpha_{+}}\right\rceil$.
@@
In what it follows we demonstrate how the single inequality given by Eq. (8), is indeed necessary and sufficient to describe the convex hull given by conditions Eq. (6) and (7), for any dimension $D$. See also Appendix A for a numerical study.
The if part of the proof follows from an exhaustive analytical characterization of the extremal points defined by the constraints Eq. (6),(7), together with the D1D-1 ordering conditions $\{\lambda_{k}\leq\lambda_{k+1}\}_{k=0,1,..,D-2}$ and the PSD constraint $\lambda_{0}\geq 0$. The feasible set of eigenvalues that satisfy these conditions forms a polytope in the $\mathbb{R}^{D-1}$ affine subspace defined by the normalization condition $\sum_{i=0}^{D-1}\lambda_{i}=1$ (see the dark shadow polytope in Figure 1). In the dual picture, such a polytope can be equivalently described as a convex combination of extremal points (vertices). Each vertex is necessarily specified by the saturation of D1D-1 inequalities and the fulfillment of the remaining ones. The proof relies on the (analytical) computation of all such points and the verification that they are detected either by Eq. (6) or Eq. (7).
To begin with the characterization of such vertex, we first consider the case $\{\lambda_{k}=\lambda_{k+1}\}_{k=0,1,\dots,D-2}$. Under normalization, it leads to $\bm{\lambda}_{\rm MMS}=\left(1/D,1/D,...,1/D\right)$, corresponding to the MMS, which is detected by both conditions. The second option is to consider Eq. (8) and pick D2D-2 conditions (i.e., all except one) from $\{\lambda_{k}=\lambda_{k+1}\}_{k=0,1,..,D-2}$. Removing the $r+1$-th condition from the last set implies that the vector is of the form $\bm{\lambda}^{(r)}=(\underbrace{a,...,a}_{r},\underbrace{b,...,b}_{D-r})$ with $a\leq b$. Notice that geometrically, $\bm{\lambda}^{(r)}$ generate a cone whose apex corresponds to the MMS which is enclosed by the D1D-1 eigenvalue ordering conditions.The normalization condition $r\cdot a+(D-r)\cdot b=1$ constraints these two quantities. Let us find $a\leq 1/D$ and $b\geq 1/D$ such that the proposed inequality Eq. (8) is also saturated by $\bm{\lambda}^{(r)}$. For $r\leq c$, we obtain $a=(r+\alpha_{-})/[r(D+\alpha_{-})],b=1/(D+\alpha_{-})$, which is detected by Eq. (7). On the other hand, for $r>c$, we obtain $a=1/(D+\alpha_{+}),b=(D-r+\alpha_{+})/[(D-r)(D+\alpha_{+})]$, which is detected by Eq. (6). The remaining cases involve $\lambda_{0}=0$ (i.e., rank-deficient states). The only solution here is for $\alpha_{-}=-1$, r=1,a=b=1/(D1)r=1,a=b=1/(D-1), since considering $r>1$ or $0\geq\alpha_{-}>-1$ leads to incompatible equations for all $D$.
The only if part is a consequence of the fact that condition Eq. (8) is Schur-convex, and it is saturated by the extremals of both simplices, i.e. $\bm{\lambda}_{\pm}=\left(\frac{1}{D+\alpha_{\pm}},\cdots,\frac{1}{D+\alpha_{\pm}},1-\frac{D-1}{D+\alpha_{\pm}}\right)$. Further elaboration on this claim can be found in the Appendix B. ∎
The previous lemma, together with the convexity of the AS set, enables to derive new sufficient conditions for AS, which are strictly stronger than those proposed in Theorem 3.
Theorem 4.
Let $\rho$ be a normalized bipartite state acting in a Hilbert space $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ of dimension $D=N\cdot M$, and let $\bm{\lambda}=\{\lambda_{i}\}_{i=0}^{D-1}$ denote its corresponding eigenvalues in non-decreasing order, i.e., $0\leq\lambda_{0}\leq\lambda_{1}\cdots\ \leq\lambda_{D-1}$ with $\sum_{i=0}^{D-1}\lambda_{i}=1$. If
$3\cdot\sum_{i=0}^{\lceil\frac{D-1}{3}\rceil-1}\lambda_{i}+\left(D+2-3\cdot\left\lceil\frac{D-1}{3}\right\rceil\right)\cdot\lambda_{\lceil\frac{D-1}{3}\rceil}\geq 1,$ (9)
then $\rho$ is AS.
Proof.
The proof follows from Lemma 1 by considering $\alpha_{+}=2$ and $\alpha_{-}=-1$ (Eq. (4)). ∎
Let us summarize our findings and illustrate them: the single inequality given by Eq. (9) is satisfied by all states detected by $\lambda_{0}\geq 1/(D+2)$ and $\lambda_{D-1}\leq 1/(D-1)$. Moreover, it is also satisfied by states not detected by the previous conditions but which belong to the convex hull of both simplexes. For example, consider a bipartite $2$-qubit state, with total dimension $D=4$, whose eigenvalues are given by $\bm{\lambda}=(1/12,1/4,1/4,5/12)$. This state cannot be certified AS through the conditions of Theorem 3 which correspond to $\lambda_{0}\geq 1/6$ nor by $\lambda_{D-1}\leq 1/3$, yet the state is detected by Eq. (9). This ensures the existence of at least one convex decomposition onto states that are detected either from $\lambda_{0}$ or $\lambda_{D-1}$. Indeed, such decomposition is given by the equally weighted contribution of the vertex of each simplex $\bm{\lambda}=\frac{1}{2}\left[(1/6,1/6,1/6,1/2)+(0,1/3,1/3,1/3)\right]$.
@@
The inequality given by Eq. (9) involve several eigenvalues, however, it can be relaxed to a hierarchy set of weaker conditions for AS, containing fewer eigenvalues. Given the non-decreasing ordering, $\{\lambda_{k}\leq\lambda_{k+1}\}$, one can replace $\lambda_{k+1}$ by $\lambda_{k}$ in Eq. (9), to yield new sufficient conditions for AS. Albeit being weaker, such conditions are based on fewer eigenvalues than the original theorem, which enables the detection of AS from incomplete knowledge of the spectrum. By iteratively applying the above substitution for $i\leq\kappa\leq\lceil\frac{D-1}{3}\rceil$, we derive the following hierarchy of AS conditions:
@@
$\kappa$ $=0:\quad 3\cdot\lambda_{0}+\cdots+3\cdot\lambda_{\lceil\frac{D-1}{3}\rceil-1}+(D+2-3\cdot\left\lceil\frac{D-1}{3}\right\rceil)\cdot\lambda_{\lceil\frac{D-1}{3}\rceil}\geq 1,$ (10)
@@
$\kappa$ $=1:\quad 3\cdot\lambda_{0}+\cdots+(D+5-3\cdot\left\lceil\frac{D-1}{3}\right\rceil)\cdot\lambda_{\lceil\frac{D-1}{3}\rceil-1}+\geq 1,$
@@
$\kappa$ $=\left\lceil\frac{D-1}{3}\right\rceil:\quad(D+2)\cdot\lambda_{0}\geq 1.$
Each $\kappa$-level in Eq. (10) is weaker than the previous $\kappa-1$, but does not require the knowledge of eigenvalue $\lambda_{\kappa}$, allowing also for some intermediate eigenvalues to be unknown. By considering the two smallest eigenvalues, we derive the weaker condition $(D-1)\lambda_{1}+3\lambda_{0}\geq 1$, which already includes the $2D$ vertices of both simplexes despite being suboptimal. The whole polytope of AS states is enclosed by all the possible permutations on the eigenvalues of inequality Eq. (9), but just as with the case of the simplexes in Figure 1, one linear inequality is enough for ordered eigenvalues (see Lemma 1).
3.2 Extending Gurvits-Barnum criterion for arbitrary dimensions
Let us observe that condition $\lambda_{\max}(\rho)\leq 1/(N\cdot M-1)$ is completely enclosed within the Gurvits-Barnum ball. Thus, we focus on the set of states detected by the convex hull (CH) of the ball (Theorem 1) plus the condition $\lambda_{\min}(\rho)\geq 1/(N\cdot M+2)$ (Theorem 3). This approach improves the existing separability criteria for the general dimension $D=N\cdot M$. Notably, the Hilbert-Schmidt distance between the vertices of the simplex and the maximally mixed state tends towards twice the radius of the ball as $N\cdot M\rightarrow\infty$. Figure 2 provides a 2D-sketch of the geometry of such a convex hull and how different regions are detected.
Figure 2: Detected absolutely separable (AS) states. Left panel: Schematic two-dimensional representation of the convex hull (orange) of the Gurvits-Barnum ball (yellow) and our simplex (blue) in barycentric coordinates (cf. Figure 1). Right panel: Purity $\mathrm{Tr}(\rho^{2})$ and minimal eigenvalue $\lambda_{\min}(\rho)$ of the detected states for a 22-qutrit system. The points in the striped region do not correspond to physical states.
While the analytical expression for the convex hull (Figure 2) in high dimensions is long and challenging to work with, it is possible to express the problem as a disciplined convex program (DCP), as formalized in the next proposition.
@@ 3.2 Extending Gurvits-Barnum criterion for arbitrary dimensions
$\begin{array}[]{crl}\min_{\{\varrho_{1}\geq 0,\varrho_{2}\geq 0\}}&0&\\ \mbox{s.t.}&\rho&=\varrho_{1}+\varrho_{2}\\ &\lambda_{\min}(\varrho_{1})&\geq\frac{\mathrm{Tr}(\varrho_{1})}{D+2}\\ &\mathrm{Tr}(\varrho^{2}_{2})&\leq\frac{\mathrm{Tr}(\varrho_{2})^{2}}{D-1}\end{array}$ (11)
@@
If the problem is feasible, it implies that the algorithm indeed succeeded in finding $\{\varrho_{1},\varrho_{2}\}$ fulfilling the constraints. Hence, both states $\{\varrho_{1},\varrho_{2}\}$ are certified AS from the positivity of the inverse map introduced in Theorem 1 or from the Gurvits-Barnum ball defined in Theorem 3 (numerically expressed as bounds on the Frobenius norm $\rho_{2}-\mbox{Tr}(\rho_{2})\cdot\mathds{1}/D||_{F}\leq\mbox{Tr}(\rho_{2})/\sqrt{D(D-1)}$) , respectively. Finally, by convexity, their sum $\rho$ is also AS. ∎
@@ 3.3 Unifying separability criteria into a single convex program
Here, we outline how to generalize the disciplined convex program from Proposition 1 into potentially stronger sufficient conditions for separability. Although such separability criteria, as already discussed, may arise from a plethora of different techniques, it is interesting to notice that all of them can be ultimately cast as a set of inequalities, $\{f_{i}(\rho)\leq 0\}$, where the functions $\{f_{i}\}$ are convex in the state $\rho$. Let us designate by $\mathcal{R}_{i}=\{\rho\geq 0\mbox{ s.t. }f_{i}(\rho)\leq 0\}$ the set of (unnormalized) states detected by each criterion $i$. From the curvature of $\{f_{i}\}$, the sets $\{\mathcal{R}_{i}\}$ are convex as well. However, such sets, stemming from different conditions, may be incomparable. How can we build a new condition that includes all of them? The following theorem offers a solution to this question and generalizes Proposition 1.
@@
Let $\rho$ be a quantum state, and $\{f_{i}\}$ a set of convex functions leading to a set of non comparable inequalities $\{f_{i}(\rho)\leq 0\}$, each of them corresponding to a sufficient condition for separability. If the convex program
[…]
@@
By construction, the previous theorem implicitly tests whether the state $\rho$ belongs to the convex hull of $\mathcal{R}$, where $\mathcal{R}=\cup_{i}\mathcal{R}_{i}$, is the union of the detected sets [38]. Consequently, Theorem 5 enables the certification of states $\rho$ that are not detected by any criteria $\{f_{i}(\rho)\leq 0\}$ alone, but that nonetheless can be decomposed into states that are detected by such criteria. In turn, convex problems of the structure Eq. (12), are not only mathematically well-posed [39, 40], but for some instances they may admit polynomial-time solutions. In this regard, there is a number of available solvers tailored to these problems [41], as well as user-friendly parsers [42].